If 3.00 moles of H2O are produced, how many grams of oxygen must be consumed?



1
Expert's answer
2020-07-20T06:22:48-0400

 3.00 moles of H2O are produced Have Mass =( 3 * 18 ) = 54 g .


 3.00 moles of H2O are produced Have 3 moles of "O" ATOM .


Mass of O Atom = ( 3 * 16 ) g = 48 g . Answer .